# Renormalization group approach to graphene bilayers

### v1

@@ III.2.1 Derivation of the RG flow: general principle
$g_{k}=\lim_{{\bf p}\rightarrow{\bf 0}}\frac{1}{a!}\frac{1}{b!}\frac{\partial^{a}}{\partial({\bf p}^{2})^{a}}\frac{\partial^{b}}{\partial(p_{c}^{2})^{b}}\Gamma^{(2)}_{k,ij}[{\bf r};{\bf p},-{\bf p}]\Big|_{{\bf r}_{0,k}},$ (27)
for appropriately chosen values of the indices $a,b,c,i,j$. In Eq. (27), ${\bf r}_{0,k}$ denotes the running flat-phase configuration, see Eq. (4), which writes here:
@@
$\partial_{t}g_{k}=\lim_{{\bf p}\rightarrow{\bf 0}}\frac{1}{a!}\frac{1}{b!}\frac{\partial^{a}}{\partial({\bf p}^{2})^{a}}\frac{\partial^{b}}{\partial(p_{c}^{2})^{b}}\Big\{\partial_{t}\Gamma^{(2)}_{k,ij}[{\bf r};{\bf p},-{\bf p}]\Big|_{{\bf r}_{0,k}}+\int_{\bf q}\Gamma^{(3)}_{k,ijl}[{\bf r};{\bf p},-{\bf p},{\bf q}]\,\partial_{t}r_{k}^{\,l}({\bf q})\Big|_{{\bf r}_{0,k}}\Big\}$ (29)
which couples the RG flow of $\Gamma^{(2)}_{k,ij}[{\bf r};{\bf p},-{\bf p}]$ to higher functional derivatives of $\Gamma_{k}[{\bf r}]$. The former is obtained by taking two functional derivatives of the Wetterich equation Eq. (19), which yields
@@
where $G_{k,ab}[{\bf r},{\bf q}]$ is the the (modified) propagator :
$G_{k,ab}[{\bf r},{\bf q}]=\big[\Gamma^{(2)}_{k,ab}[{\bf r};{\bf q},-{\bf q}]+R_{k,ab}({\bf q})\big]^{-1}.$ (31)
@@
Note, however, that due to translational invariance any configuration of the form Eq. (33), with $\zeta_{k}$ replaced by an arbitrary constant $\theta_{k}$, would also satisfy Eq. (32). To lift this ambiguity, it is more convenient to define $\zeta_{k}$ through a stationarity condition with respect to the tangent vector of the membrane, namely
@@ III.2.2 Propagators and spectrum of excitations
$\Gamma_{k,p}^{(2)}[{\bf r};\mathbf{q},-\mathbf{q}]\Big|_{{\bf r}_{0,k}}\hskip-11.38092pt=\begin{pmatrix}Z_{k}\mathbf{q}^{4}+M_{1}(\mathbf{q}^{2})&\zeta_{k}^{2}(\lambda_{k}+\mu_{k})\,q_{1}q_{2}\\[2.0pt] \zeta_{k}^{2}(\lambda_{k}+\mu_{k})\,q_{1}q_{2}&Z_{k}\mathbf{q}^{4}+M_{2}(\mathbf{q}^{2})\end{pmatrix}$, (37)
@@
$M_{i}(\mathbf{q}^{2})=\zeta_{k}^{2}\big(\mu_{k}\,\mathbf{q}^{2}+(\lambda_{k}+\mu_{k})\,q_{i}^{2}\big),$ (38)
@@
$\Gamma^{(2)}_{k}[{\bf r};\mathbf{q},-\mathbf{q}]+R_{k}({\bf q})$ (40)
@@
$\Gamma_{k,ij}^{(2)}[{\bf r};\mathbf{q},-\mathbf{q}]\Big|_{{\bf r}_{0,k}}+R_{k}({\bf q})$ $=G_{k,1}^{-1}({\bf q})\,P^{\perp}_{ij}(\boldsymbol{q})$ (44)
@@
The expression Eq. (44) is easily inversed according to the properties $(P^{\parallel})^{2}=(P^{\perp})^{2}=1$ and $P^{\parallel}P^{\perp}=0$ and one gets:
$\left[\Gamma_{k,ij}^{(2)}[{\bf r};\mathbf{q},-\mathbf{q}]\big|_{{\bf r}_{0,k}}+R_{k,ij}({\bf q})\right]^{-1}$ $=G_{k,1}({\bf q})P^{\perp}_{ij}(\boldsymbol{q})$ (45)
@@
    $G_{k,1}(\boldsymbol{q})=\frac{1}{Z_{k}\,\boldsymbol{q}^{4}+m_{1,k}^{2}\,\boldsymbol{q}^{2}+R_{k}(\boldsymbol{q})}\,;$ (46)
@@
    $G_{k,2}(\boldsymbol{q})=\frac{1}{Z_{k}\,\boldsymbol{q}^{4}+m_{2,k}^{2}\,\boldsymbol{q}^{2}+R_{k}(\boldsymbol{q})}\,;$ (47)
@@
    $G_{k,0}(\boldsymbol{q})=\frac{1}{Z_{k}\,\boldsymbol{q}^{4}+R_{k}(\boldsymbol{q})}\,.$ (48)
@@ III.2.3 RG equations
$\left\{\begin{aligned} \mu_{k}&=\frac{1}{\zeta_{k}^{2}}\lim_{\boldsymbol{p}\to 0}\frac{\partial}{\partial\boldsymbol{p}^{2}}\,\Gamma_{k,22}^{(2)}[{\bf r};\boldsymbol{p},-\boldsymbol{p}]\Big|_{{\bf r}_{0,k}},\\[2.84544pt] \lambda_{k}&=\frac{1}{\zeta_{k}^{2}}\lim_{\boldsymbol{p}\to 0}\frac{\partial}{\partial p_{2}^{2}}\,\Gamma_{k,22}^{(2)}[{\bf r};\boldsymbol{p},-\boldsymbol{p}]\Big|_{{\bf r}_{0,k}}-\mu_{k},\\[2.84544pt] Z_{k}&=\lim_{{\bf p}\to 0}\,\frac{\partial^{2}}{\partial({\bf p}^{2})^{2}}\,\Gamma^{(2)}_{k,33}[{\bf r};{\bf p},-{\bf p}]\Big|_{{\bf r}_{0,k}},\end{aligned}\right$. (49)
@@
$\hskip 22.76228pt+\frac{1}{\zeta_{k}^{2}}\int_{\bf q}\Gamma_{k,22j}^{(3)}[{\bf r};\boldsymbol{p},-\boldsymbol{p},\boldsymbol{q}]\,\partial_{t}r_{j}(\boldsymbol{q})\bigg]\Bigg|_{{\bf r}={\bf r}_{0,k}}\hskip-22.76228pt,$
@@
(with no summation over the index $i$), one obtains the RG flow equations for the running couplings $\lambda_{k}$, $\mu_{k}$, $\zeta_{k}$ and for the field renormalization $Z_{k}$.
@@
where the propagators are given by
$G_{k,i}(\boldsymbol{q})^{-1}=P(\boldsymbol{q})+m_{i,k}^{2}\,\boldsymbol{q}^{2},\qquad P(\boldsymbol{q})=Z_{k}\,\boldsymbol{q}^{4}+R_{k}(\boldsymbol{q}),$ (58)
with
$m_{0,k}^{2}=0,\qquad m_{1,k}^{2}=\zeta_{k}^{2}\,\mu_{k},\qquad m_{2,k}^{2}=\zeta_{k}^{2}\,(\lambda_{k}+2\,\mu_{k})\,.$ (59)
@@
Their explicit expressions in terms of dimensionless quantities are given in App. B.1.
@@ III.2.4 RG equations in the flat phase
$\left\{\begin{aligned} &L_{a00}^{2+\alpha}=-\frac{1}{4\,A_{2}}\,\widehat{\partial}_{t}\int_{q}\boldsymbol{q}^{\alpha}\,P(\boldsymbol{q})^{-a}\\ &N_{a00}^{2+\alpha}=-\frac{1}{4\,A_{2}}\,\widehat{\partial}_{t}\int_{q}\boldsymbol{q}^{\alpha}\,\frac{\partial P(\boldsymbol{q})}{\partial\boldsymbol{q}^{2}}\,P(\boldsymbol{q})^{-a}\ .\end{aligned}\right$. (66)
The dimensionless threshold functions $\bar{N}_{a00}^{2+\alpha}$ that enter in Eq. (65) are related to the dimensionful threshold function by:
$N_{abc}^{2+\alpha}=(Z_{k}k^{4})^{-a-b-c}Z_{k}k^{4+\alpha}\bar{N}_{abc}^{2+\alpha}\ .$ (67)
Their expressions in terms of dimensionless quantities are given in Appendix B.1.
@@ III.3 Bilayers
$\hskip-8.5359pt+\frac{g_{1k}}{8\ell_{k}^{4}}(\boldsymbol{s}^{2}-\ell_{k}^{2})^{2}+\frac{g_{2k}}{2\ell_{k}^{2}}(\boldsymbol{s}.\partial_{\alpha}\boldsymbol{r})^{2}+\frac{g_{3k}}{4\ell_{k}^{2}}\big({\bf s}^{2}-\ell_{k}^{2})u_{\alpha\alpha}\bigg\}.$
Starting from this expression, we will perform a few simplifications. First, we will consider the case of an infinite coupling $g_{1k}$, so that the constraint $|{\bf S}|^{2}=\ell_{k}^{2}$ is effectively enforced. It then follows automatically that the volume-area term $({\bf S}^{2}-\ell_{k}^{2})\,u_{\alpha\alpha}$, proportional to $g_{3k}$, does not contribute, so that one puts $g_{3k}\equiv 0$. Next, in the flat phase, we consider both a low-wavector limit, ${\bf q}^{2}\ell^{2}\ll\zeta^{2}$ and a regime where $g_{2k}$ is very large which simplifies the spectrum of excitations. Owing to the fact that these two limits do not commute, this procedure leads to a very specific form of the flexural propagator and, ultimately, to the mechanical rigidity crossover, see below.
@@ III.3.2 Propagators
$\Gamma_{k,rr}^{(2)}[{\bf r},{\bf s};\mathbf{q},-\mathbf{q}]|_{FP}=\begin{pmatrix}2Z_{k}{\bf q}^{4}+2M_{1}[{\bf q}^{2}]&2\zeta_{k}^{2}(\lambda_{k}+\mu_{k})q_{1}q_{2}\\ \\ 2\zeta_{k}^{2}(\lambda_{k}+\mu_{k})q_{1}q_{2}&2Z_{k}{\bf q}^{4}+2M_{2}[{\bf q}^{2}]\end{pmatrix}$ (75)
$\Gamma_{k,rs}^{(2)}[{\bf r},{\bf s};\mathbf{q},-\mathbf{q}]|_{FP}=\begin{pmatrix}2Z_{k}{\bf q}^{4}+q_{1}^{2}g_{2k}&-\frac{ig_{2k}\zeta_{k}q_{1}}{\ell_{k}}&\frac{ig_{2k}\zeta_{k}q_{1}}{\ell_{k}}\\ \frac{ig_{2k}\zeta_{k}q_{1}}{\ell_{k}}&\frac{1}{2}Z_{k}{\bf q}^{4}+\frac{1}{2}\zeta_{k}^{2}\frac{2g_{2k}}{\ell_{k}^{2}}+\frac{1}{2}M_{1}[{\bf q}^{2}]&\frac{1}{2}\zeta_{k}^{2}(\lambda_{k}+\mu_{k})q_{1}q_{2}\\ \frac{ig_{2k}\zeta_{k}q_{2}}{\ell_{k}}&\frac{1}{2}\zeta_{k}^{2}(\lambda_{k}+\mu_{k})q_{1}q_{2}&\frac{1}{2}Z_{k}{\bf q}^{4}+\frac{1}{2}\zeta_{k}^{2}\frac{2g_{2k}}{\ell_{k}^{2}}+\frac{1}{2}M_{2}[{\bf q}^{2}]\end{pmatrix}$ (76)
with $M_{i}[{\bf q}^{2}]=\zeta_{k}^{2}\big(\mu_{k}{\bf q}^{2}+(\lambda_{k}+\mu_{k})q_{i}^{2}\big)$ and
@@
The eigenvalues of $\left[\Gamma_{k,ij}^{(2)}|_{FP}+R_{k}(\mathbf{q})\right]^{-1}$ can be easily extracted.
@@
$\left\{\begin{array}[]{ll}\displaystyle G_{k,1rr}(\boldsymbol{q})=\frac{1}{2\big(Z_{k}\,\boldsymbol{q}^{4}+m_{1,k}^{2}\boldsymbol{q}^{2}+R_{k}(\boldsymbol{q})\big)}\\ \\ \displaystyle G_{k,2rr}(\boldsymbol{q})=\frac{1}{2\big(Z_{k}\,\boldsymbol{q}^{4}+m_{2,k}^{2}\boldsymbol{q}^{2}+R_{k}(\boldsymbol{q})\big)}\ .\end{array}\right$. (78)
@@
$G_{k,ss}(\boldsymbol{q})=\frac{2}{Z_{k}\,\boldsymbol{q}^{4}+2m_{ss,k}^{2}+R_{k}(\boldsymbol{q})}\,.$ (79)
@@
$G_{k,1rs}(\boldsymbol{q})=\frac{2}{Z_{k}\,\boldsymbol{q}^{4}+m_{1,k}^{2}\boldsymbol{q}^{2}+m_{rs,k}^{2}+R_{k}(\boldsymbol{q})}\,$, (81)
@@
Since we are interested in long-distance physics, we may consider the low-wavevector limit, ${\bf q}^{2}\,\ell^{2}\ll\zeta^{2}$ together with that of large-$g_{2}$ coupling constant. In this regime, a second propagator takes the form
$G_{k,2rs}(\boldsymbol{q})=\frac{2}{Z_{k}\,\boldsymbol{q}^{4}+m_{2,k}^{2}\boldsymbol{q}^{2}+m_{rs,k}^{2}+R_{k}(\boldsymbol{q})}\,.$ (83)
@@
The second, and most important, propagator obtained in the low-wavevector, large-$g_{2}$ limits reads
$G_{k,3rs}(\boldsymbol{q})=\frac{1}{2\big((Z_{k}+c_{k})\,\boldsymbol{q}^{4}+R_{k}(\boldsymbol{q})\big)}\,$, (84)
@@
The propagator $G_{k,3rs}(\boldsymbol{q})$ is associated with the flexural mode of the mean membrane and carries neither a mass nor a pseudo-mass. As can be seen from Eq. (84), its explicit dependence on $g_{2,k}$ has disappeared (see App. C). However, because the limits $\boldsymbol{q}\to 0$ and $g_{2,k}\to 0$ do not commute, the signature of a finite interlayer coupling $g_{2,k}$ remains encoded in the specific form of $G_{k,3rs}$.

### v2

@@ III.2.1 Derivation of the RG flow: general principle
$g_{k,ij}=\lim_{{\bf p}\rightarrow{\bf 0}}\frac{1}{a!}\frac{1}{b!}\frac{\partial^{a}}{\partial({\bf p}^{2})^{a}}\frac{\partial^{b}}{\partial(p_{c}^{2})^{b}}\Gamma^{(2)}_{k,ij}[{\bf r};{\bf p},-{\bf p}]\Big|_{{\bf r}_{0,k}},$ (27)
– without summation over $a,b$ indices – for appropriately chosen values of the indices $a,b,c,i,j$. In Eq. (27), ${\bf r}_{0,k}$ denotes the running flat-phase configuration, see Eq. (4), which writes here:
@@
$\partial_{t}g_{k,ij}=\lim_{{\bf p}\rightarrow{\bf 0}}\frac{1}{a!}\frac{1}{b!}\frac{\partial^{a}}{\partial({\bf p}^{2})^{a}}\frac{\partial^{b}}{\partial(p_{c}^{2})^{b}}\Big\{\partial_{t}\Gamma^{(2)}_{k,ij}[{\bf r};{\bf p},-{\bf p}]\Big|_{{\bf r}_{0,k}}+\int_{\bf q}\Gamma^{(3)}_{k,ijl}[{\bf r};{\bf p},-{\bf p},{\bf q}]\,\partial_{t}r_{l}({\bf q})\Big|_{{\bf r}_{0,k}}\Big\}$ (29)
– without summation over $a,b$ indices – where $r_{l}$ denotes the $l$-th component of the position field $\bf r$}. Eq.(29) couples the RG flow of $\Gamma^{(2)}_{k,ij}[{\bf r};{\bf p},-{\bf p}]$ to higher functional derivatives of $\Gamma_{k}[{\bf r}]$. The former is obtained by taking two functional derivatives of the Wetterich equation Eq. (19), which yields
@@
where $G_{k,ab}[{\bf r},{\bf q}]$ is the (modified) propagator :
$G_{k,ab}[{\bf r},{\bf q}]=\big[\Gamma^{(2)}_{k,ab}[{\bf r};{\bf q},-{\bf q}]+{\cal R}_{k,ab}({\bf q})\big]^{-1}.$ (31)
@@
Note, however, that because of translational invariance any configuration of the form Eq. (33), with $\zeta_{k}$ replaced by an arbitrary constant $\theta_{k}$, would also satisfy Eq. (32). To lift this ambiguity, it is more convenient to define $\zeta_{k}$ through a stationarity condition with respect to the tangent vector of the membrane, namely
@@ III.2.2 Propagators and spectrum of excitations
$\Gamma_{k,p}^{(2)}[{\bf r};\mathbf{q},-\mathbf{q}]\Big|_{{\bf r}_{0,k}}\hskip-11.38092pt=\begin{pmatrix}Z_{k}\mathbf{q}^{4}+M_{1}(\mathbf{q})&\zeta_{k}^{2}(\lambda_{k}+\mu_{k})\,q_{1}q_{2}\\[2.0pt] \zeta_{k}^{2}(\lambda_{k}+\mu_{k})\,q_{1}q_{2}&Z_{k}\mathbf{q}^{4}+M_{2}(\mathbf{q})\end{pmatrix}$, (37)
@@
$M_{i}(\mathbf{q})\equiv\zeta_{k}^{2}\big(\mu_{k}\,\mathbf{q}^{2}+(\lambda_{k}+\mu_{k})\,q_{i}^{2}\big),$ (38)
@@
$\Gamma^{(2)}_{k}[{\bf r};\mathbf{q},-\mathbf{q}]+{\cal R}_{k}({\bf q})$ (40)
@@
$\Gamma_{k,ij}^{(2)}[{\bf r};\mathbf{q},-\mathbf{q}]\Big|_{{\bf r}_{0,k}}+{\cal R}_{k}({\bf q})$ $=G_{k,1}^{-1}({\bf q})\,P^{\perp}_{ij}(\boldsymbol{q})$ (44)
@@
The expression Eq. (44) is easily inversed according to the properties $(P^{\parallel})^{2}=(P^{\perp})^{2}=\mbox{l\hskip-5.50003pt1}$ and $P^{\parallel}P^{\perp}=0$ and one gets:
$\left[\Gamma_{k,ij}^{(2)}[{\bf r};\mathbf{q},-\mathbf{q}]\big|_{{\bf r}_{0,k}}+{\cal R}_{k,ij}({\bf q})\right]^{-1}$ $=G_{k,1}({\bf q})P^{\perp}_{ij}(\boldsymbol{q})$ (45)
@@
    $G_{k,1}(\boldsymbol{q})=\frac{1}{Z_{k}\,\boldsymbol{q}^{4}+m_{1,k}^{2}\,\boldsymbol{q}^{2}+{\cal R}_{k}(\boldsymbol{q})}\,;$ (46)
@@
    $G_{k,2}(\boldsymbol{q})=\frac{1}{Z_{k}\,\boldsymbol{q}^{4}+m_{2,k}^{2}\,\boldsymbol{q}^{2}+{\cal R}_{k}(\boldsymbol{q})}\,;$ (47)
@@
    $G_{k,0}(\boldsymbol{q})=\frac{1}{Z_{k}\,\boldsymbol{q}^{4}+{\cal R}_{k}(\boldsymbol{q})}\,.$ (48)
@@ III.2.3 RG equations
$\left\{\begin{aligned} \mu_{k}&=\frac{1}{\zeta_{k}^{2}}\lim_{\boldsymbol{p}\to 0}\frac{\partial}{\partial\boldsymbol{p}^{2}}\ \Gamma_{k,22}^{(2)}[{\bf r};\boldsymbol{p},-\boldsymbol{p}]\Big|_{{\bf r}_{0,k}}\\[2.84544pt] \lambda_{k}&=\frac{1}{\zeta_{k}^{2}}\lim_{\boldsymbol{p}\to 0}\frac{\partial}{\partial p_{2}^{2}}\,\Gamma_{k,22}^{(2)}[{\bf r};\boldsymbol{p},-\boldsymbol{p}]\Big|_{{\bf r}_{0,k}}-\mu_{k}\\[2.84544pt] Z_{k}&=\lim_{{\bf p}\to 0}\,\frac{\partial^{2}}{\partial({\bf p}^{2})^{2}}\,\Gamma^{(2)}_{k,33}[{\bf r};{\bf p},-{\bf p}]\Big|_{{\bf r}_{0,k}}\end{aligned}\right$. (49)
@@
$\hskip 22.76228pt+\frac{1}{\zeta_{k}^{2}}\int_{\bf q}\Gamma_{k,22j}^{(3)}[{\bf r};\boldsymbol{p},-\boldsymbol{p},\boldsymbol{q}]\,\partial_{t}r_{j}(\boldsymbol{q})\bigg]\Bigg|_{{\bf r}={\bf r}_{0,k}}\hskip-22.76228pt$
@@
– without summation over the indices $i$ and $j$ – one obtains the RG flow equations for the running couplings $\lambda_{k}$, $\mu_{k}$, $\zeta_{k}$ and for the field renormalization $Z_{k}$.
@@
where the propagators $G_{k,i}(\boldsymbol{q})$, $i=0\dots 2$ are given by Eqs.(46-48).
@@
– without summation over the indices $a$, $b$, $c$ and $\alpha$. Their explicit expressions in terms of dimensionless quantities are given in App. B.1.
@@ III.2.4 RG equations in the flat phase
$\left\{\begin{aligned} &L_{a00}^{2+\alpha}=-\frac{1}{4\,A_{2}}\,\widehat{\partial}_{t}\int_{q}\boldsymbol{q}^{\alpha}\,P(\boldsymbol{q})^{-a}\\ &N_{a00}^{2+\alpha}=-\frac{1}{4\,A_{2}}\,\widehat{\partial}_{t}\int_{q}\boldsymbol{q}^{\alpha}\,\frac{\partial P(\boldsymbol{q})}{\partial\boldsymbol{q}^{2}}\,P(\boldsymbol{q})^{-a}\end{aligned}\right$. (64)
with $P(\boldsymbol{q})=Z_{k}\,\boldsymbol{q}^{4}+{\cal R}_{k}(\boldsymbol{q})$. The dimensionless threshold functions $\bar{N}_{a00}^{2+\alpha}$ that enter in Eq. (63) are related to the dimensionful threshold function by:
$N_{abc}^{2+\alpha}=(Z_{k}k^{4})^{-a-b-c}Z_{k}k^{4+\alpha}\bar{N}_{abc}^{2+\alpha}$ (65)
– without summation over the indices $a$, $b$, $c$ and $\alpha$. Their expressions in terms of dimensionless quantities are given in Appendix B.1.
@@ III.3 Bilayers
$\hskip-8.5359pt+\frac{g_{1k}}{8\ell_{k}^{4}}(\boldsymbol{s}^{2}-\ell_{k}^{2})^{2}+\frac{g_{2k}}{2\ell_{k}^{2}}(\boldsymbol{s}.\partial_{\alpha}\boldsymbol{r})^{2}+\frac{g_{3k}}{4\ell_{k}^{2}}\big({\bf s}^{2}-\ell_{k}^{2})u_{\alpha\alpha}\bigg\}$
where $u_{\alpha\beta}=1/2(\partial_{\alpha}{\bf r}.\partial_{\beta}{\bf r}-\partial_{\alpha}{\bf r}_{0}.\partial_{\beta}{\bf r}_{0})$.
Starting from this expression, we will perform a few simplifications. First, we will consider the case of an infinite coupling $g_{1k}$, so that the constraint $|{\bf S}|^{2}=\ell_{k}^{2}$ is effectively enforced. It then follows automatically that the volume-area term $({\bf S}^{2}-\ell_{k}^{2})\,u_{\alpha\alpha}$, proportional to $g_{3k}$, does not contribute, so that one puts $g_{3k}\equiv 0$. Next, in the flat phase, we consider both a low-wavevector limit, ${\bf q}^{2}\ell^{2}\ll\zeta^{2}$ and a regime where $g_{2k}$ is very large which simplifies the spectrum of excitations. Owing to the fact that these two limits do not commute, this procedure leads to a very specific form of the flexural propagator and, ultimately, to the mechanical rigidity crossover, see below.
@@ III.3.2 Propagators
$\Gamma_{k,rr}^{(2)}[{\bf r},{\bf s};\mathbf{q},-\mathbf{q}]|_{FP}=\begin{pmatrix}2Z_{k}{\bf q}^{4}+2M_{1}({\bf q})&2\zeta_{k}^{2}(\lambda_{k}+\mu_{k})q_{1}q_{2}\\ \\ 2\zeta_{k}^{2}(\lambda_{k}+\mu_{k})q_{1}q_{2}&2Z_{k}{\bf q}^{4}+2M_{2}({\bf q})\end{pmatrix}$ (73)
$\Gamma_{k,rs}^{(2)}[{\bf r},{\bf s};\mathbf{q},-\mathbf{q}]|_{FP}=\begin{pmatrix}2Z_{k}{\bf q}^{4}+q_{1}^{2}g_{2k}&-\frac{ig_{2k}\zeta_{k}q_{1}}{\ell_{k}}&\frac{ig_{2k}\zeta_{k}q_{1}}{\ell_{k}}\\ \frac{ig_{2k}\zeta_{k}q_{1}}{\ell_{k}}&\frac{1}{2}Z_{k}{\bf q}^{4}+\frac{1}{2}\zeta_{k}^{2}\frac{2g_{2k}}{\ell_{k}^{2}}+\frac{1}{2}M_{1}({\bf q})&\frac{1}{2}\zeta_{k}^{2}(\lambda_{k}+\mu_{k})q_{1}q_{2}\\ \frac{ig_{2k}\zeta_{k}q_{2}}{\ell_{k}}&\frac{1}{2}\zeta_{k}^{2}(\lambda_{k}+\mu_{k})q_{1}q_{2}&\frac{1}{2}Z_{k}{\bf q}^{4}+\frac{1}{2}\zeta_{k}^{2}\frac{2g_{2k}}{\ell_{k}^{2}}+\frac{1}{2}M_{2}({\bf q})\end{pmatrix}$ (74)
with $M_{i}({\bf q})=\zeta_{k}^{2}\big(\mu_{k}{\bf q}^{2}+(\lambda_{k}+\mu_{k})q_{i}^{2}\big)$ and
@@
The eigenvalues of $\left[\Gamma_{k,ij}^{(2)}|_{FP}+{\cal R}_{k}(\mathbf{q})\right]^{-1}$ can be easily extracted.
@@
$\left\{\begin{array}[]{ll}\displaystyle G_{k,1rr}(\boldsymbol{q})=\frac{1}{2\big(Z_{k}\,\boldsymbol{q}^{4}+m_{1,k}^{2}\boldsymbol{q}^{2}+{\cal R}_{k}(\boldsymbol{q})\big)}\\ \\ \displaystyle G_{k,2rr}(\boldsymbol{q})=\frac{1}{2\big(Z_{k}\,\boldsymbol{q}^{4}+m_{2,k}^{2}\boldsymbol{q}^{2}+{\cal R}_{k}(\boldsymbol{q})\big)}\ .\end{array}\right$. (76)
@@
$G_{k,ss}(\boldsymbol{q})=\frac{2}{Z_{k}\,\boldsymbol{q}^{4}+2m_{ss,k}^{2}+{\cal R}_{k}(\boldsymbol{q})}\,.$ (77)
@@
$G_{k,1rs}(\boldsymbol{q})=\frac{2}{Z_{k}\,\boldsymbol{q}^{4}+m_{1,k}^{2}\boldsymbol{q}^{2}+m_{rs,k}^{2}+{\cal R}_{k}(\boldsymbol{q})}\,$, (79)
@@
Since we are interested in long-distance physics, we may consider the low-wavevector limit, ${\bf q}^{2}\,\ell_{k}^{2}\ll\zeta_{k}^{2}$ at leading order in the parameter $u=q^{2}l_{k}^{2}/\zeta_{k}^{2}$. In this regime, a second propagator takes the form
$G_{k,2rs}(\boldsymbol{q})=\frac{2}{Z_{k}\,\boldsymbol{q}^{4}+m_{2,k}^{2}\boldsymbol{q}^{2}+m_{rs,k}^{2}+{\cal R}_{k}(\boldsymbol{q})}\,.$ (81)
@@
The third, and most important, propagator in the $rs$ sector, in the low-wavevector limit, reads
$G_{k,3rs}(\boldsymbol{q})=\frac{1}{2\big((Z_{k}+c_{k})\,\boldsymbol{q}^{4}+{\cal R}_{k}(\boldsymbol{q})\big)}\,$, (82)
@@
The propagator $G_{k,3rs}(\boldsymbol{q})$ is associated with the flexural mode of the mean membrane and carries neither a mass nor a pseudo-mass. As can be seen from Eq. (82), its explicit dependence on $g_{2,k}$ has disappeared (see App. C). However, as the low-wavevector expansion, $\boldsymbol{q}^{2}\ell_{k}^{2}\ll\zeta_{k}^{2}$, is not uniform in $g_{2k}$, the limits $\boldsymbol{q}\to 0$ and $g_{2,k}\to 0$ do not commute so that the signature of a finite interlayer coupling $g_{2,k}$ remains encoded in the specific form of $G_{k,3rs}$; see App. D for details.